# Cherry picking in forests: A new characterization for the unrooted hybrid number of two phylogenetic trees

### v3

@@ Abstract.
Key words and phrases:
phylogenetic network; forest; TBR distance; cherry picking sequence; hybrid number
1991 Mathematics Subject Classification:
05C05; 92D15
The second author thanks the New Zealand Marsden Fund for their financial support.
1. Introduction
Phylogenetic networks are a special type of graph which generalize phylogenetic trees and that are used to model non-treelike evolutionary processes such as recombination and hybridization [11]. There are several classes of phylogenetic networks (see e.g. [20] for a recent review), but in this paper, we shall restrict our attention to unrooted phylogenetic networks (see e.g. [13]). Such a network is a simple, connected graph ${\mathcal{N}}=(V,E)$ with leaf set $X$, where $X$ is some set of species, and every internal vertex in ${\mathcal{N}}$ has degree three. The reticulation number $r({\mathcal{N}})$ of ${\mathcal{N}}$ is defined as $|E|-(|V|-1)$. If $r({\mathcal{N}})=0$, then ${\mathcal{N}}$ has no cycles, in which case it is called an (unrooted binary) phylogenetic tree on $X$. An example of a phylogenetic network ${\mathcal{N}}$ with leaf set $X=\{1,\dots,6\}$ and $r({\mathcal{N}})=1$ is given in the left of Figure 1 (the tree in the centre of this figure is a phylogenetic tree); formal definitions are given in the next section.
One approach commonly used to construct an unrooted phylogenetic network is to take as input a set ${\mathcal{P}}$ of phylogenetic trees and to infer a network that embeds or ‘displays’ [13, 16] each tree in ${\mathcal{P}}$ (see Figure 1 for an example of a phylogenetic tree displayed in a network). Such a network always exists since, for example one can be constructed using the so-called display graph [5] for all trees in ${\mathcal{P}}$ and subsequently refining the vertices of that graph whose degree is greater than three by introducing new edges. From a biological viewpoint, it is of interest to not reconstruct an arbitrary unrooted phylogenetic network that displays each element in ${\mathcal{P}}$, but one whose reticulation number is minimized over all such networks. This is known as the Unrooted Hybridization Number problem  [13]. If ${\mathcal{N}}$ is an unrooted phylogenetic network that displays each tree in ${\mathcal{P}}$ and whose reticulation number is minimized over all such networks, then $r({\mathcal{N}})$ is denoted by $h({\mathcal{P}})$ and is called the hybrid number of ${\mathcal{P}}$.
@@ 1. Introduction
Figure 1. Left: A phylogenetic network ${\mathcal{N}}$ with leaf set $X=\{1,\dots,6\}$. Middle: A forest ${\mathcal{F}}$ that contains a single phylogenetic tree. Right: A forest ${\mathcal{F}}^{\prime}$ comprising two components. Both ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ are displayed by ${\mathcal{N}}$; the way in which ${\mathcal{F}}^{\prime}$ is displayed by ${\mathcal{N}}$ is indicated in bold.
@@
Intriguingly, in [13, Theorem 3] it is shown that, if ${\mathcal{P}}$ contains only two phylogenetic trees ${\mathcal{T}}$ and ${\mathcal{T}}^{\prime}$ on the same leaf set, then $h({\mathcal{P}})$ is in fact equal to the tree bisection and reconnection (TBR) distance between ${\mathcal{T}}$ and ${\mathcal{T}}^{\prime}$ (see e.g. [1] and also [25]). Informally, this distance is defined as the minimum number of so-called tree bisection and reconnection moves required to convert ${\mathcal{T}}$ into ${\mathcal{T}}^{\prime}$, where such a move on a phylogenetic tree is essentially the process of cutting an edge in that tree and then reconnecting the resulting two trees by introducing a new edge between two edges, one from each of the two trees. The TBR distance has been studied for several years, and is still a topic of current research [6, 17, 18, 24, 26]. Furthermore, computing this distance is an NP-hard optimization problem [1]. Hence, gaining a better understanding of the hybrid number of two phylogenetic trees is not only of interest for constructing phylogenetic networks but also for shedding new light on computing the TBR distance between them.
@@
In this paper, we establish a new characterization for the hybrid number of two phylogenetic trees (Theorem 4.1) and, therefore the TBR distance (Theorem 4.2). Our approach is motivated by the concept of a cherry picking sequence that was introduced in [10] in the context of two rooted (binary) phylogenetic trees. A cherry in such a tree is a pair of leaves that are adjacent to the same vertex. Roughly speaking, for a pair of rooted phylogenetic trees ${\mathcal{R}}$ and ${\mathcal{R}}^{\prime}$ on the same leaf set $X$, a cherry picking sequence for ${\mathcal{R}}$ and ${\mathcal{R}}^{\prime}$ is an ordering or sequence $\sigma$ of the elements in $X$, say ${\color[rgb]{0,0,0}\sigma}=(x_{1},x_{2},\ldots,x_{n})$, $n=|X|$, such that each $x_{i}$, $i\in\{1,2,\ldots,n\}$, is a leaf in a cherry in the phylogenetic trees obtained from ${\mathcal{R}}$ and ${\mathcal{R}}^{\prime}$ by pruning off $x_{1},x_{2},\ldots,x_{i-1}$. In [10] it is shown that a cherry picking sequence for two rooted phylogenetic trees ${\mathcal{R}}$ and ${\mathcal{R}}^{\prime}$ exists precisely if ${\mathcal{R}}$ and ${\mathcal{R}}^{\prime}$ can be embedded in a rooted binary phylogenetic network ${\mathcal{N}}$ on $X$ that satisfies certain structural properties (namely it is ‘time consistent’ and ‘tree-child’). The number of elements $x_{i}$ in $\sigma$ with $i\in\{1,2,\ldots,n\}$ for which the two trees obtained from ${\mathcal{R}}$ and ${\mathcal{R}}^{\prime}$ by pruning off $x_{1},x_{2},\ldots,x_{i-1}$ have cherries $\{x_{i},y_{i}\}$ and $\{x_{i},y_{i}^{\prime}\}$ such that $y_{i}\neq y_{i}^{\prime}$ is called the weight of $\sigma$. In addition, it is shown in the same paper that the minimum weight over all cherry picking sequences for ${\mathcal{R}}$ and ${\mathcal{R}}^{\prime}$ equates to the minimum number of reticulations (vertices with in-degree two and out-degree one) that are required to display ${\mathcal{R}}$ and ${\mathcal{R}}^{\prime}$ by a rooted phylogenetic network that is time consistent and tree-child. Since the publication of [10], cherry picking sequences have been extensively studied and generalized to larger classes of rooted phylogenetic networks and arbitrarily large collections of rooted phylogenetic trees (see e.g. [15, 23]). This theoretical work has, in turn, resulted in the development of practical algorithms to reconstructed rooted phylogenetic networks from a set of rooted phylogenetic trees [4, 12]. In a related line of research, cherry picking operations have recently been used to define and compute distances between phylogenetic networks [21, 22].
@@
In this paper, we shall introduce an analog of a cherry picking sequence for two unrooted phylogenetic trees ${\mathcal{T}}$ and ${\mathcal{T}}^{\prime}$ on the same leaf set. In particular, the main result of this paper (Theorem 4.1) proves that the hybrid number and, consequently, the TBR distance of ${\mathcal{T}}$ and ${\mathcal{T}}^{\prime}$ can be given in terms of the minimum weight of a cherry picking sequence using an alternative weight to the rooted setting, and the minimum is again taken over all such sequences for ${\mathcal{T}}$ and ${\mathcal{T}}^{\prime}$. Interestingly, to obtain this result we found that it is necessary to define a cherry picking sequence for two forests on the same leaf set (a forest is a collection of phylogenetic trees, see e.g. the right-hand side of Figure 1), and then apply such a sequence to the special case where the two forests are both phylogenetic trees. Intuitively, this is the case because, instead of only pruning leaves of cherries as in the rooted case that is described in the previous paragraph, a cherry picking sequence whose weight characterizes the hybrid number of ${\mathcal{T}}$ and ${\mathcal{T}}^{\prime}$ allows for operations that split a phylogenetic tree into two smaller phylogenetic trees by deleting edges that are not necessarily pendant. It is also interesting to note that our definition of a cherry picking sequence has some similarity with certain data reduction rules that have been recently introduced for establishing parameterized algorithms to compute the TBR distance between two phylogenetic trees [18]. We now summarize the content of the rest of this paper. In Section 2, we present some preliminaries. Subsequently, in Section 3, we introduce the main concept of a cherry picking sequence for two forests, and show that such a sequence always exists for any pair of forests. In Section 4, we show that if $\sigma$ is a cherry picking sequence for two forests ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ on $X$, then there is an unrooted phylogenetic network ${\mathcal{N}}$ on $X$ that displays ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ such that $r({\mathcal{N}})$ is at most the weight of $\sigma$ (Theorem 4.5). In Section 5, we prove four technical lemmas that are concerned with how forests are displayed in networks. These lemmas allow us to complete the proof of our main result (Theorem 4.1) in Section 6, where we prove that if ${\mathcal{N}}$ is a phylogenetic network that displays two forests ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ on $X$, then there is a cherry picking sequence for ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ whose weight is at most $r({\mathcal{N}})$ (Theorem 6.1). We conclude in Section 7 with a brief discussion of some future directions of research.
2. Preliminaries
@@
Graphs. A graph $G=(V,E)$ is an ordered pair consisting of a set of vertices $V=V(G)$ and a (multi-)set of edges $E=E(G)$. All graphs $G$ considered in this paper are undirected multi-graphs, that is, they do not have loops but might contain vertices $u$ and $v$ such that there is more than one edge in $G$ joining $u$ and $v$. We call such edges of $G$ multi-edges. In case there is a single edge in $G$ joining two vertices $u$ and $v$ we denote it by $\{u,v\}$; if there is at most one edge between any pair of vertices, then $E(G)$ is a set. An edge $e$ of a graph $G$ is called a cut-edge of $G$ if the deletion of $e$ disconnects $G$. If $e$ is a cut-edge that contains a vertex of $G$ of degree one then we call $e$ a trivial cut-edge. A subdivision of $G$ is a graph obtained from $G$ by replacing edges in $G$ by paths containing at least one edge. If $G$ has two edges $\{u,v\}$, $\{v,w\}$, $u,v,w\in V(G)$ and $v$ has degree two and is not incident with two edges $\{u,v\}$ or two edges $\{v,w\}$, the process of replacing these two edges by a single edge $\{u,w\}$ and removing $v$ is called suppressing the vertex $v$. If $G$ has a multi-edge between two vertices $u$ and $v$ in $V(G)$, the process of replacing the edges by a single edge $\{u,v\}$ is called suppressing the multi-edge. A leaf of $G$ is a vertex in $V(G)$ of degree at most one. The set of leaves of $G$ is denoted by $L(G)$. Finally, we define $r(G)=|E|-(|V|-1)$ and refer to $r(G)$ as the reticulation number of $G$ which is also known as the cyclomatic number of $G$ [3]. Phylogenetic trees and networks. Throughout the paper, $X$ denotes a finite set with $n=|X|\geq 1$, where we think of $X$ as a set of species.
@@
An edge of ${\mathcal{N}}$ that contains a leaf $x$ of ${\mathcal{N}}$ is called a pendant edge in ${\mathcal{N}}$ (so that, in particular, $x$ is a proper pendant subtree of ${\mathcal{N}}$ and $e_{x}$ is the pendant edge of $x$). In case ${\mathcal{N}}$ has at least two leaves, a cherry of ${\mathcal{N}}$ consists of two distinct leaves $x$ and $y$ of ${\mathcal{N}}$ such that either ${\mathcal{N}}$ is equal to the edge $\{x,y\}$, or $x$ and $y$ are adjacent to a common vertex (so that, in particular, in the latter case the phylogenetic tree $\{x,y\}$ is a proper pendant subtree of ${\mathcal{N}}$). We shall denote a cherry consisting of leaves $x$ and $y$ by $(x,y)$, where the order of $x$ and $y$ is unimportant. We call ${\mathcal{N}}$ pendantless if it contains no proper pendant subtree with at least two leaves (e.g. for the phylogenetic network ${\mathcal{N}}$ in Figure 1, since the subtree with leaf set $\{5,6\}$ is a cherry of ${\mathcal{N}}$, it follows that ${\mathcal{N}}$ is not pendantless). We also need notation for two further special cases. Suppose that ${\mathcal{T}}$ is a pendant subtree of ${\mathcal{N}}$. If ${\mathcal{T}}$ has leaf set $\{x,y,z\}\subseteq X$ and ${\mathcal{T}}$ is a proper pendant subtree of ${\mathcal{N}}$ such that $(x,y)$ is a cherry of ${\mathcal{N}}$, then we denote ${\mathcal{T}}$ by $((x,y),z)$ (for brevity, if ${\mathcal{T}}={\mathcal{N}}$, then we shall also denote ${\mathcal{T}}$ by $((x,y),z)$ although the choice of the cherry $(x,y)$ is unimportant). In addition, if ${\mathcal{T}}$ has leaf set $\{x,y,z,w\}\subseteq X$ and $(x,y)$, $(z,w)$ are both cherries in ${\mathcal{N}}$, then we denote ${\mathcal{T}}$ by $((x,y),(z,w))$.
@@
For an edge $e$ in a phylogenetic tree ${\mathcal{T}}$ on $X$ we denote by ${\mathcal{T}}-e$ the forest obtained from ${\mathcal{T}}$ by deleting $e$ and suppressing degree-two vertices (if any) in the resulting trees. If ${\mathcal{F}}$ is a forest on $X$, $|X|\geq 2$, and $x\in X$, then we let ${\mathcal{T}}_{x}$ denote the tree in ${\mathcal{F}}$ that contains $x$ in its leaf set. We denote by ${\mathcal{F}}-x$ the forest obtained from ${\mathcal{F}}$ by removing ${\mathcal{T}}_{x}$ from ${\mathcal{F}}$ in case $|L({\mathcal{T}}_{x})|=1$ and replacing ${\mathcal{T}}_{x}$ by ${\mathcal{T}}_{x}-x$ otherwise. Also, for $e$ an edge in ${\mathcal{F}}$, we let ${\mathcal{T}}_{e}$ denote the tree in ${\mathcal{F}}$ with $e$ in its edge set, and denote by ${\mathcal{F}}-e$ the forest obtained by replacing ${\mathcal{T}}_{e}$ by ${\mathcal{T}}_{e}-e$. For example, referring to Figure 1 for $x=4$, the tree ${\mathcal{T}}_{x}$ in ${\mathcal{F}}^{\prime}$ is $((2,4),(5,6))$ and ${\mathcal{F}}^{\prime}-4$ is the forest comprising the cherry with leaf set $\{1,3\}$ and the three-leaf tree $((5,6),2)$. Similarly, for $e$ being the pendant edge incident with $4$, the tree ${\mathcal{T}}_{e}$ in ${\mathcal{F}}^{\prime}$ is again $((2,4),(5,6))$ and ${\mathcal{F}}-e$ is the forest comprising the cherry with leaf set $\{1,3\}$, the isolated vertex 4, and the three-leaf tree $((5,6),2)$.
@@
Displaying phylogenetic trees and forests. Suppose that ${\mathcal{N}}$ is a phylogenetic network on $X$, and that ${\mathcal{T}}$ is a phylogenetic tree on some subset $Y\subseteq X$. Then we say that ${\mathcal{T}}$ is displayed by ${\mathcal{N}}$ if ${\mathcal{T}}$ can be obtained from a subtree ${\mathcal{N}}[{\mathcal{T}}]$ of ${\mathcal{N}}$ by suppressing all vertices in ${\mathcal{N}}[{\mathcal{T}}]$ with degree two. We shall refer to ${\mathcal{N}}[{\mathcal{T}}]$ as an image of ${\mathcal{T}}$ in ${\mathcal{N}}$. For example in Figure  1, for ${\mathcal{T}}=((2,4),(5,6))$, a subgraph ${\mathcal{N}}[{\mathcal{T}}]$ of ${\mathcal{N}}$ with leaf set $\{2,4,5,6\}$ is indicated in bold. Note that an image of ${\mathcal{T}}$ in ${\mathcal{N}}$ is isomorphic to a subdivision of ${\mathcal{T}}$ and that ${\mathcal{T}}$ could have several images in ${\mathcal{N}}$. Usually the image of ${\mathcal{T}}$ in ${\mathcal{N}}$ that we are considering is clear from the context, but in case we want to make this clearer we shall refer to it explicitly.

### v5

@@ Abstract
keywords:
phylogenetic network, forest, TBR distance, cherry picking sequence, hybrid number
1 Introduction
Phylogenetic networks are a special type of graph which generalize phylogenetic trees and that are used to model non-treelike evolutionary processes such as recombination and hybridization [12]. There are several classes of phylogenetic networks (see e.g. [21] for a recent review), but in this paper, we shall restrict our attention to unrooted phylogenetic networks (see e.g. [14]). Such a network is a simple, connected graph ${\mathcal{N}}=(V,E)$ with leaf set $X$, where $X$ is some set of species, and every internal vertex in ${\mathcal{N}}$ has degree three. The reticulation number $r({\mathcal{N}})$ of ${\mathcal{N}}$ is defined as $|E|-(|V|-1)$. If $r({\mathcal{N}})=0$, then ${\mathcal{N}}$ has no cycles, in which case it is called an (unrooted binary) phylogenetic tree on $X$. An example of a phylogenetic network ${\mathcal{N}}$ with leaf set $X=\{1,\dots,6\}$ and $r({\mathcal{N}})=1$ is given in the left of Fig. 1 (the tree in the centre of this figure is a phylogenetic tree); formal definitions are given in the next section.
One approach commonly used to construct an unrooted phylogenetic network is to take as input a set ${\mathcal{P}}$ of phylogenetic trees and to infer a network that embeds or ‘displays’ [14, 17] each tree in ${\mathcal{P}}$ (see Fig. 1 for an example of a phylogenetic tree displayed in a network). Such a network always exists since, for example one can be constructed using the so-called display graph [5] for all trees in ${\mathcal{P}}$ and subsequently refining the vertices of that graph whose degree is greater than three by introducing new edges. From a biological viewpoint, it is of interest to not reconstruct an arbitrary unrooted phylogenetic network that displays each element in ${\mathcal{P}}$, but one whose reticulation number is minimized over all such networks. This is known as the Unrooted Hybridization Number problem  [14]. If ${\mathcal{N}}$ is an unrooted phylogenetic network that displays each tree in ${\mathcal{P}}$ and whose reticulation number is minimized over all such networks, then $r({\mathcal{N}})$ is denoted by $h({\mathcal{P}})$ and is called the hybrid number of ${\mathcal{P}}$.
@@ 1 Introduction
Figure 1: Left: A phylogenetic network ${\mathcal{N}}$ with leaf set $X=\{1,\dots,6\}$. Middle: A forest ${\mathcal{F}}$ that contains a single phylogenetic tree. Right: A forest ${\mathcal{F}}^{\prime}$ comprising two components. Both ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ are displayed by ${\mathcal{N}}$; the way in which ${\mathcal{F}}^{\prime}$ is displayed by ${\mathcal{N}}$ is indicated in bold.
@@
Intriguingly, in [14, Theorem 3] it is shown that, if ${\mathcal{P}}$ contains only two phylogenetic trees ${\mathcal{T}}$ and ${\mathcal{T}}^{\prime}$ on the same leaf set, then $h({\mathcal{P}})$ is in fact equal to the tree bisection and reconnection (TBR) distance between ${\mathcal{T}}$ and ${\mathcal{T}}^{\prime}$ (see e.g. [1] and also [26]). Informally, this distance is defined as the minimum number of so-called tree bisection and reconnection moves required to convert ${\mathcal{T}}$ into ${\mathcal{T}}^{\prime}$, where such a move on a phylogenetic tree is essentially the process of cutting an edge in that tree and then reconnecting the resulting two trees by introducing a new edge between two edges, one from each of the two trees. The TBR distance has been studied for several years, and is still a topic of current research [6, 18, 19, 25, 27]. Furthermore, computing this distance is an NP-hard optimization problem [1]. Hence, gaining a better understanding of the hybrid number of two phylogenetic trees is not only of interest for constructing phylogenetic networks but also for shedding new light on computing the TBR distance between them.
@@
In this paper, we establish a new characterization for the hybrid number of two phylogenetic trees (Theorem 4.1) and, therefore the TBR distance (Theorem 4.2). Our approach is motivated by the concept of a cherry picking sequence that was introduced in [11] in the context of two rooted (binary) phylogenetic trees. A cherry in such a tree is a pair of leaves that are adjacent to the same vertex. Roughly speaking, for a pair of rooted phylogenetic trees ${\mathcal{R}}$ and ${\mathcal{R}}^{\prime}$ on the same leaf set $X$, a cherry picking sequence for ${\mathcal{R}}$ and ${\mathcal{R}}^{\prime}$ is an ordering or sequence $\sigma$ of the elements in $X$, say ${\color[rgb]{0,0,0}\sigma}=(x_{1},x_{2},\ldots,x_{n})$, $n=|X|$, such that each $x_{i}$, $i\in\{1,2,\ldots,n\}$, is a leaf in a cherry in the phylogenetic trees obtained from ${\mathcal{R}}$ and ${\mathcal{R}}^{\prime}$ by pruning off $x_{1},x_{2},\ldots,x_{i-1}$. In [11] it is shown that a cherry picking sequence for two rooted phylogenetic trees ${\mathcal{R}}$ and ${\mathcal{R}}^{\prime}$ exists precisely if ${\mathcal{R}}$ and ${\mathcal{R}}^{\prime}$ can be embedded in a rooted binary phylogenetic network ${\mathcal{N}}$ on $X$ that satisfies certain structural properties (namely it is ‘time consistent’ and ‘tree-child’). The number of elements $x_{i}$ in $\sigma$ with $i\in\{1,2,\ldots,n\}$ for which the two trees obtained from ${\mathcal{R}}$ and ${\mathcal{R}}^{\prime}$ by pruning off $x_{1},x_{2},\ldots,x_{i-1}$ have cherries $\{x_{i},y_{i}\}$ and $\{x_{i},y_{i}^{\prime}\}$ such that $y_{i}\neq y_{i}^{\prime}$ is called the weight of $\sigma$. In addition, it is shown in the same paper that the minimum weight over all cherry picking sequences for ${\mathcal{R}}$ and ${\mathcal{R}}^{\prime}$ equates to the minimum number of reticulations (vertices with in-degree two and out-degree one) that are required to display ${\mathcal{R}}$ and ${\mathcal{R}}^{\prime}$ by a rooted phylogenetic network that is time consistent and tree-child. Since the publication of [11], cherry picking sequences have been extensively studied and generalized to larger classes of rooted phylogenetic networks and arbitrarily large collections of rooted phylogenetic trees (see e.g. [16, 24]). This theoretical work has, in turn, resulted in the development of practical algorithms to reconstruct rooted phylogenetic networks from a set of rooted phylogenetic trees [4, 13]. In a related line of research, cherry picking operations have recently been used to define and compute distances between phylogenetic networks [22, 23].
@@
In this paper, we shall introduce an analog of a cherry picking sequence for two unrooted phylogenetic trees ${\mathcal{T}}$ and ${\mathcal{T}}^{\prime}$ on the same leaf set. In particular, the main result of this paper (Theorem 4.1) proves that the hybrid number and, consequently, the TBR distance of ${\mathcal{T}}$ and ${\mathcal{T}}^{\prime}$ can be given in terms of the minimum weight of a cherry picking sequence using an alternative weight to the rooted setting, and the minimum is again taken over all such sequences for ${\mathcal{T}}$ and ${\mathcal{T}}^{\prime}$. Interestingly, to obtain this result we found that it is necessary to define a cherry picking sequence for two forests on the same leaf set (a forest is a collection of phylogenetic trees, see e.g. the right-hand side of Fig. 1), and then apply such a sequence to the special case where the two forests are both phylogenetic trees. Intuitively, this is the case because, instead of only pruning leaves of cherries as in the rooted case that is described in the previous paragraph, a cherry picking sequence whose weight characterizes the hybrid number of ${\mathcal{T}}$ and ${\mathcal{T}}^{\prime}$ allows for operations that split a phylogenetic tree into two smaller phylogenetic trees by deleting edges that are not necessarily pendant. It is also interesting to note that our definition of a cherry picking sequence has some similarity with certain data reduction rules that have been recently introduced for establishing parameterized algorithms to compute the TBR distance between two phylogenetic trees [19].
We now summarize the content of the rest of this paper. In Section 2, we present some preliminaries. Subsequently, in Section 3, we introduce the main concept of a cherry picking sequence for two forests, and show that such a sequence always exists for any pair of forests. In Section 4, we show that if $\sigma$ is a cherry picking sequence for two forests ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ on $X$, then there is an unrooted phylogenetic network ${\mathcal{N}}$ on $X$ that displays ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ such that $r({\mathcal{N}})$ is at most the weight of $\sigma$ (Theorem 4.7). In Section 5, we prove four technical lemmas that are concerned with how forests are displayed in networks. These lemmas allow us to complete the proof of our main result (Theorem 4.1) in Section 6, where we prove that if ${\mathcal{N}}$ is a phylogenetic network that displays two forests ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ on $X$, then there is a cherry picking sequence for ${\mathcal{F}}$ and ${\mathcal{F}}^{\prime}$ whose weight is at most $r({\mathcal{N}})$ (Theorem 6.1). We conclude in Section 7 with a brief discussion of some future directions of research.
2 Preliminaries
@@
Graphs. The graphs that we shall consider in this paper will be multi-graphs. We denote such a graph $G$ by an ordered pair $G=(V,E)$, where $V=V(G)$ are the vertices of $G$ and $E=E(G)$ are the edges of $G$. No graph considered will contain a loop, but it may contain a multi-edge, that is, a pair of vertices which are joined by more than one edge. In case there is a single edge in $G$ joining two vertices $u$ and $v$ we call it an edge and denote it by $\{u,v\}$. Note that if $E(G)$ contains only edges, then $E(G)$ is a set. An edge $e$ of a graph $G$ is called a cut-edge of $G$ if the deletion of $e$ disconnects $G$. If $e$ is a cut-edge that contains a vertex of $G$ of degree one then we call $e$ a trivial cut-edge. A subdivision of $G$ is a graph obtained from $G$ by replacing edges in $G$ by paths containing at least one edge. If $G$ has two edges $\{u,v\}$, $\{v,w\}$, $u,v,w\in V(G)$ and $v$ has degree two and is not incident with two edges $\{u,v\}$ or two edges $\{v,w\}$, the process of replacing these two edges by a single edge $\{u,w\}$ and removing $v$ is called suppressing the vertex $v$. If $G$ has a multi-edge between two vertices $u$ and $v$ in $V(G)$, the process of replacing the edges by a single edge $\{u,v\}$ is called suppressing the multi-edge. A leaf of $G$ is a vertex in $V(G)$ of degree at most one. The set of leaves of $G$ is denoted by $L(G)$. Finally, we define $r(G)=|E|-(|V|-1)$ and refer to $r(G)$ as the reticulation number of $G$ which is also known as the cyclomatic number of $G$ [3].
Phylogenetic trees and networks. Throughout the paper, $X$ denotes a finite set with $n=|X|\geq 1$, where we think of $X$ as a set of species.
@@
An edge of ${\mathcal{N}}$ that contains a leaf $x$ of ${\mathcal{N}}$ is called a pendant edge in ${\mathcal{N}}$ (so that, in particular, $x$ is a proper pendant subtree of ${\mathcal{N}}$ and $e_{x}$ is the pendant edge of $x$). In case ${\mathcal{N}}$ has at least two leaves, a cherry of ${\mathcal{N}}$ consists of two distinct leaves $x$ and $y$ of ${\mathcal{N}}$ such that either ${\mathcal{N}}$ is equal to the edge $\{x,y\}$, or $x$ and $y$ are adjacent to a common vertex (so that, in particular, in the latter case the phylogenetic tree $\{x,y\}$ is a proper pendant subtree of ${\mathcal{N}}$). We shall denote a cherry consisting of leaves $x$ and $y$ by $(x,y)$, where the order of $x$ and $y$ is unimportant. We call ${\mathcal{N}}$ pendantless if it contains no proper pendant subtree with at least two leaves (e.g. for the phylogenetic network ${\mathcal{N}}$ in Fig. 1, since the subtree with leaf set $\{5,6\}$ is a cherry of ${\mathcal{N}}$, it follows that ${\mathcal{N}}$ is not pendantless). We also need notation for two further special cases. Suppose that ${\mathcal{T}}$ is a pendant subtree of ${\mathcal{N}}$. If ${\mathcal{T}}$ has leaf set $\{x,y,z\}\subseteq X$ and ${\mathcal{T}}$ is a proper pendant subtree of ${\mathcal{N}}$ such that $(x,y)$ is a cherry of ${\mathcal{N}}$, then we denote ${\mathcal{T}}$ by $((x,y),z)$ (for brevity, if ${\mathcal{T}}={\mathcal{N}}$, then we shall also denote ${\mathcal{T}}$ by $((x,y),z)$ although the choice of the cherry $(x,y)$ is unimportant). In addition, if ${\mathcal{T}}$ has leaf set $\{x,y,z,w\}\subseteq X$ and $(x,y)$, $(z,w)$ are both cherries in ${\mathcal{N}}$, then we denote ${\mathcal{T}}$ by $((x,y),(z,w))$.
@@
For an edge $e$ in a phylogenetic tree ${\mathcal{T}}$ on $X$ we denote by ${\mathcal{T}}-e$ the forest obtained from ${\mathcal{T}}$ by deleting $e$ and suppressing degree-two vertices (if any) in the resulting trees. If ${\mathcal{F}}$ is a forest on $X$, $|X|\geq 2$, and $x\in X$, then we let ${\mathcal{T}}_{x}$ denote the tree in ${\mathcal{F}}$ that contains $x$ in its leaf set. We denote by ${\mathcal{F}}-x$ the forest obtained from ${\mathcal{F}}$ by removing ${\mathcal{T}}_{x}$ from ${\mathcal{F}}$ in case $|L({\mathcal{T}}_{x})|=1$ and replacing ${\mathcal{T}}_{x}$ by ${\mathcal{T}}_{x}-x$ otherwise. Also, for $e$ an edge in ${\mathcal{F}}$, we let ${\mathcal{T}}_{e}$ denote the tree in ${\mathcal{F}}$ with $e$ in its edge set, and denote by ${\mathcal{F}}-e$ the forest obtained by replacing ${\mathcal{T}}_{e}$ by ${\mathcal{T}}_{e}-e$. For example, referring to Fig. 1 for $x=4$, the tree ${\mathcal{T}}_{x}$ in ${\mathcal{F}}^{\prime}$ is $((2,4),(5,6))$ and ${\mathcal{F}}^{\prime}-4$ is the forest comprising the cherry with leaf set $\{1,3\}$ and the three-leaf tree $((5,6),2)$. Similarly, for $e$ being the pendant edge incident with $4$, the tree ${\mathcal{T}}_{e}$ in ${\mathcal{F}}^{\prime}$ is again $((2,4),(5,6))$ and ${\mathcal{F}}-e$ is the forest comprising the cherry with leaf set $\{1,3\}$, the isolated vertex 4, and the three-leaf tree $((5,6),2)$.
@@
Displaying phylogenetic trees and forests. Suppose that ${\mathcal{N}}$ is a phylogenetic network on $X$, and that ${\mathcal{T}}$ is a phylogenetic tree on some subset $Y\subseteq X$. Then we say that ${\mathcal{T}}$ is displayed by ${\mathcal{N}}$ if ${\mathcal{T}}$ can be obtained from a subtree ${\mathcal{N}}[{\mathcal{T}}]$ of ${\mathcal{N}}$ by suppressing all vertices in ${\mathcal{N}}[{\mathcal{T}}]$ with degree two. We shall refer to ${\mathcal{N}}[{\mathcal{T}}]$ as an image of ${\mathcal{T}}$ in ${\mathcal{N}}$. For example in Fig.  1, for ${\mathcal{T}}=((2,4),(5,6))$, a subgraph ${\mathcal{N}}[{\mathcal{T}}]$ of ${\mathcal{N}}$ with leaf set $\{2,4,5,6\}$ is indicated in bold. Note that an image of ${\mathcal{T}}$ in ${\mathcal{N}}$ is isomorphic to a subdivision of ${\mathcal{T}}$ and that ${\mathcal{T}}$ could have several images in ${\mathcal{N}}$. Usually the image of ${\mathcal{T}}$ in ${\mathcal{N}}$ that we are considering is clear from the context, but in case we want to make this clearer we shall refer to it explicitly.